Circular No. 7379

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000P IN NGC 4965 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000P (cf. IAUC 7378), taken by M. Calkins on Mar. 10.4 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova at an early epoch. The spectrum exhibits a very blue continuum, with P-Cyg Balmer lines and He I. The Hα and Hβ profiles show narrow emission (FWHM < 500 km/s), but this appears to be from a superimposed H II region, rather than being intrinsic to the supernova. The narrow lines provide a recession velocity of 2230 km/s for the host galaxy, and the Hβ P-Cyg profile yields an expansion velocity for the supernova of 7400 km/s.

Corrigendum. On IAUC 7378, last line, for 09^{s}.88, 13′59′′.3. read 10^{s}.56, 14′02′′.5.

AX J1747.1–2845 AND THE GALACTIC CENTER REGION H. Murakami, M. Sakano, and K. Koyama, Kyoto University; T. Fu-

rusho, Tokyo Metropolitan University; and T. Sakamoto, RIKEN, write: “A new transient x-ray source, designated AX J1747.1–2845, was discovered by on-going ultra-deep observations of the Galactic-center region with the ASCA satellite made since Mar. 11.0 UT. The position of the new transient is α = 17^{h}47^{m}07^{s}.0, δ = −28^{o}45′09′′ (equinox 2000.0; estimated error radius 1′.2, 90-percent confidence). No x-ray source has so far been reported within this error circle. The x-ray spectrum is well represented by a power law with photon index ∼ 2.5, undergoing a photoelectric absorption with a hydrogen column density of 1.0 × 10^{23} atoms/cm^{2}. The observed 2–10-keV flux is 2.3 × 10^{−10} erg s^{−1} cm^{−2}. Another transient source is detected at α = 17^{h}47^{m}01^{s}.3, δ = −28^{o}52′52′′ (error radius 1′), with a 2–10-keV flux of 2.1 × 10^{−9} erg s^{−1} cm^{−2}; the position is consistent with the x-ray burster SAX J1747.0–2853, which was recently reported to become luminous again (IAUC 7377). We have also detected the persistent source 1E 1743.1–2843 with a 2–10-keV flux of 1.0 × 10^{−10} erg s^{−1} cm^{−2}. The Galactic nuclei Sgr A^{∗} is not resolved into a point source. The 2–10-keV flux within the 3’-radius region centered on Sgr A^{∗} is 6 × 10^{−11} erg s^{−1} cm^{−2}. The ASCA observation continues until Mar. 22.0. Simultaneous observations are encouraged.”

2000 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green